# Modeling Multi-Task Model Merging as Adaptive Projective Gradient Descent

### v2

@@ Abstract
Merging multiple expert models offers a promising approach for performing multi-task learning without accessing their original data. Existing methods attempt to alleviate task conflicts by sparsifying task vectors or promoting orthogonality among them. However, they overlook the fundamental requirement of model merging: ensuring the merged model performs comparably to task-specific models on respective tasks. We find these methods inevitably discard task-specific information that, while causing conflicts, is crucial for performance. Based on our findings, we frame model merging as a constrained optimization problem (i.e., minimizing the gap between the merged model and individual models, subject to the constraint of retaining shared knowledge) and solve it via adaptive projective gradient descent. Specifically, we align the merged model with individual models by decomposing and reconstituting the loss function, alleviating conflicts through data-free optimization of task vectors. To retain shared knowledge, we optimize this objective by projecting gradients within a shared subspace spanning all tasks. Moreover, we view merging coefficients as adaptive learning rates and propose a task-aware, training-free strategy. Experiments show that our plug-and-play approach consistently outperforms previous methods, achieving state-of-the-art results across diverse architectures and tasks in both vision and NLP domains. Our code is available here.
@@ 1 Introduction
Fine-tuning on pre-trained foundational models to address downstream tasks has become an effective paradigm (Muqeeth et al., 2024). However, the independent deployment of multiple fine-tuned models increases storage costs. While traditional multi-task learning (MTL) can mitigate these issues, they typically require concurrent training across multiple task-specific datasets, which incurs significant training overhead and potential privacy risks (Wei et al., 2024). Consequently, there is a growing interest in merging multiple expert models into a unified model without accessing their original data (Yang et al., 2024a; Huang et al., 2024). Model merging is performed directly at the parameter level and maintains only one final model during inference. In recent years, numerous pre-trained and fine-tuned checkpoints have been released on open-source communities like GitHub or Hugging Face, making it easy to obtain expert models from diverse domains. These rich model repositories underscore the value of model merging.
One popular approach, Task Arithmetic (TA) (Ilharco et al., 2023), combines task vectors through arithmetic operations for model merging. A major challenge is addressing conflicts that emerge when multiple task-specific models coexist within a single model. Ties-Merging (Yadav et al., 2024) proposes pruning redundant parameters, resolving sign conflicts, and merging sparse models, while AdaMerging (Yang et al., 2024c) applies test-time adaptation techniques to adjust merging coefficients in the weight space. Most recently, AWD (Xiong et al., 2024) finds that orthogonality among task vectors is key to model merging and introduce adaptive weight disentanglement to improve orthogonality. However, these methods overlook the fundamental requirement of model merging: ensuring the merged model performs comparably to task-specific models on their respective tasks.
@@
Based on our rethinking, we model model merging as a constrained optimization problem (i.e., minimizing the gap between the merged model and individual models, subject to the constraint of retaining shared knowledge) and solve it via an adaptive projective gradient descent (doGe) method. Specifically, we measure the gap between the merged model and individual models in task-specific losses, and decompose it into a data-free objective using the first-order Taylor expansion. To alleviate conflicts, we introduce a modification vector $\Delta$ (i.e., redundant parameters) to each task vector. This data-free objective aims to achieve optimal average performance across multiple tasks by optimizing $\Delta$. For the modification vector, task vectors still compete to minimize the loss on their own tasks. Therefore, we construct a shared subspace based on all task vectors and optimize the problem within this subspace. The gradient of $\Delta$ can be divided into two components: one projected onto the shared subspace and the other orthogonal to it. We only take gradient steps in the direction orthogonal to the shared space, effectively constraining task vector optimization. As the former represents movements of parameters within the shared subspace, and the latter maintains shared knowledge while minimizing the gap for each task. Moreover, we determine task-aware, training-free merging coefficients based on the norm of task vectors to mitigate the dominance of any single task’s gradient influence.
We conduct experiments on diverse vision and NLP tasks, including classification and generation, using various fully fine-tuned and LoRA fine-tuned architectures. Our plug-and-play approach achieves up to 11.6% gains over TA, 5.8% over AdaMerging. Simple task-aware $\lambda$ provides a 2.8% performance boost. Furthermore, experiments on unseen tasks and out-of-distribution test sets demonstrate its generalization and robustness. Extensive ablation studies clarify the mechanisms of each component.
@@ 2 Related Work
Model merging. Model merging (Crisostomi et al., 2024; Wang et al., 2024) eliminates the need for raw training data or expensive computations. It operates directly at the parameter level and consolidates multiple models into a single final model for inference. Existing model merging methods are categorized into two paradigms: pre-merging and during-merging (Yang et al., 2024a). Pre-merging methods aim to create favorable conditions for merging, such as using linearized fine-tuning to achieve weight disentanglement (Ortiz-Jimenez et al., 2023; Tang et al., 2024b).
During-merging focuses on designing techniques to merge given models, and is further divided into data-free methods and test-time adaptation (TTA) methods. Data-free methods rely solely on pre-trained and fine-tuned models for weight merging (Choi et al., 2024). For example, Ties-Merging (Yadav et al., 2024) trims redundant parameters based on magnitudes, resolves parameter sign conflicts, and merges sparse models. Concrete Merging (Tang et al., 2023) introduces a meta-learning framework to learn a concrete mask that suppresses conflicting parameters. Calculating the loss gap has been reflected in some studies: MetaGPT (Zhou et al., 2024) formally defines the loss difference, deriving a closed-form solution for $\lambda$. TATR (Anonymous, 2024) defines knowledge conflict between losses and expands it as the product of gradients and task vectors. We focus on optimizing task vectors via gradient descent while constraining it within a shared subspace to retain shared knowledge.
@@
TTA methods assume access to unlabeled test datasets, which can be considered a form of transductive learning. AdaMerging (Yang et al., 2024c) learns the merging coefficients by minimizing entropy as a surrogate loss on test data. Representation Surgery (Yang et al., 2024b) calibrates biases and aligns the representation of the merged model with that of independent models. In contrast, we design a data-free objective to resolve conflicts.
@@
Multi-task learning. Existing MTL research addresses the issue of negative transfer (Jiang et al., 2023) from two principal perspectives: architecture and optimization. From the architectural perspective, negative transfer is mitigated through strategies like modularization (Lu et al., 2024), sparsification (Sun et al., 2020), or soft sharing of the backbone. From an optimization perspective, it is widely recognized that tasks sharing similar underlying structures can benefit from being trained together. Gradient alignment methods (Yu et al., 2020; Shi et al., 2022) emphasize maintaining consistency in gradient directions or signs to resolve conflicts, which projects one task’s gradient onto the normal plane of another task’s gradient to reduce forgetting (Saha et al., 2021). Our approach enhances multi-task performance by aligning the merged model with each individual model and utilizing adaptive merging coefficients.
@@ 3.1 Preliminary
We begin with a pre-trained model $f$, parameterized by $\boldsymbol{\theta}_{0}$, which has been trained on a large-scale dataset. This model is paired with a collection of $n$ downstream tasks, denoted as $\{\mathcal{D}_{i}\}_{i=1}^{n}$. Subsequently, the pre-trained model $f$ is fine-tuned individually for each downstream task $\mathcal{D}_{i}$, resulting in a series of fine-tuned models each with its unique parameters $\boldsymbol{\theta}_{i}$. To isolate task-specific information, we define the task vector as $\boldsymbol{\tau}_{i}=\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{0}$, a concept introduced by Ilharco et al. (2023). The set of these task vectors is represented as $\{\boldsymbol{\tau}_{i}\}_{i=1}^{n}$, enabling a focused analysis of the task-specific characteristics. Model merging aims to compose a multi-task model $\boldsymbol{\theta}^{*}$ to approximate the optimal solution:
@@ 3.2 Rethinking Model Merging for MTL
Resolving conflicts among tasks is a key challenge in model merging. Unlike MTL, which can mitigate conflicts during training with access to original data, model merging operates entirely in the parameter space. Existing methods mainly address conflicts by sparsely adjusting parameters, either by dropping conflicting parameters based on signs (Yadav et al., 2024) or importance scores (Du et al., 2024). Other methods promote orthogonality among task vectors, either by fine-tuning models in the tangent space (Ortiz-Jimenez et al., 2023) or directly optimizing task vectors (Xiong et al., 2024). While these methods alleviate conflicts, they inevitably discard task-specific information that contributes to conflicts, resulting in performance degradation. The fundamental requirement of model merging is that the merged model should perform comparably to task-specific models on their respective tasks—an aspect existing methods fail to consider. As shown in Fig. 1, increasing the task numbers leads to a continuous performance decline across methods. This is because more tasks result in increased negative transfer, causing the discard of valuable conflict-related task-specific knowledge. Therefore, we propose explicitly modeling the gap between the merged model $\boldsymbol{\theta}^{*}$ and individual models $\boldsymbol{\theta}_{i}$. This transforms conflict resolution into an optimization problem that can be solved using gradient descent.
@@
Figure 3: An illustration of the element magnitudes in the task vector. Best viewed when zoomed in.
@@
A critical observation is the importance of the merging coefficient $\lambda$. In methods like TA, a unified $\lambda$ is searched on the validation set. Ideally, $\lambda$ values should be task- and layer-specific. However, when dealing with a large number of tasks and layers, traditional methods such as grid search or combinatorial optimization search (Liu et al., 2020) become impractical. TTA methods require training $\lambda$ using unlabeled test data, which also presents limitations. A statistical analysis of task vector values reveals that tasks and layers exhibit different magnitudes (see Fig. 3). Modern adaptive optimizers (e.g., Adam) dynamically adjust learning rates based on gradient history, which is often more effective than using a global learning rate. These optimizers suppress parameters with large gradients and reward those with small gradients, smoothing gradient fluctuations. Similarly, task vectors $\boldsymbol{\tau}_{i}$ represent cumulative gradients for each task, and $\lambda$ can be viewed as a learning rate balancing gradients across multiple tasks.
@@ 4 Methodology
$\min_{\boldsymbol{\theta}^{*}}$ $\mathcal{L}(\Delta;\lambda_{i},\boldsymbol{\tau}_{i}):=\sum_{i=1}^{n}\|\boldsymbol{\theta}^{*}-\boldsymbol{\theta}_{i}\|_{\text{Gap}},$ (2)
$\,\mathrm{s.t.}$ $\|\boldsymbol{\theta}^{*}-(\boldsymbol{\theta}_{0}+\lambda\sum_{i=1}^{n}\boldsymbol{\tau}_{i})\|_{S_{\text{share}}}=0.$
Then, we solve it via adaptive projective gradient descent:
@@ 4.1 A Data-Free ̵Objective
Considering the fundamental requirement that the merged model should perform comparably to its respective task-specific model for each task, we follow Zhou et al. (2024) to define the objective for resolving model merging as:
@@
Similarly, calculating the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{j}(\boldsymbol{\theta}_{0})$ of the pre-trained model for task $j$ requires unavailable data $\mathcal{D}_{j}$. As an alternative, we approximate this gradient with the task vector $-\boldsymbol{\tau}_{j}$, as the task vector can be interpreted as an accumulation of gradients. Consequently, the data-free objective can be estimated as:
@@ 4.2 Shared Subspace Optimization
We utilize Singular Value Decomposition (SVD) to extract the rank-$k$ subspace for the task vector. Specifically, the first $k$ singular vectors from the left singular matrix are selected as $\boldsymbol{B}_{j}$, forming an orthogonal basis that efficiently captures the primary information within the task-specific $\boldsymbol{\tau}_{j}$. Once the subspaces for all tasks are established, they are combined into a shared subspace $S_{\text{share}}=span\{[\boldsymbol{B}_{1},...,\boldsymbol{B}_{n}]\}$. However, $S_{\text{share}}$ includes overlapping singular vectors, indicating redundant parameters in the weight space across tasks. Such overlaps challenge the orthogonality requirement of basis vectors and lead to inaccuracies during projection onto the shared subspace. To mitigate this, we perform another SVD on $S_{\text{share}}$ to further deduplicate it, resulting in a refined $S_{\text{share}}$ that effectively preserves shared knowledge.
@@
Compared to optimizing $\Delta$ in the original parameter space, our approach explicitly constrains the gradient directions that the optimizer can take. By taking gradient steps in the direction orthogonal to the shared subspace, we narrow the gap with the task-specific model. This effectively mitigates task conflicts while retaining shared knowledge.
4.3 Task-aware Training-free $\lambda$
@@
       Perform SVD on $\boldsymbol{\tau}_{i}$: $\boldsymbol{\tau}_{i}=U_{i}\Sigma_{i}V_{i}^{\top}$
@@
$S_{\text{share}}\leftarrow$ the first $k$ columns of concatenated $[\boldsymbol{B}_{1},\dots,\boldsymbol{B}_{n}]$

### v3

@@ Abstract
Merging multiple expert models offers a promising approach for performing multi-task learning without accessing their original data. Existing methods attempt to alleviate task conflicts by sparsifying task vectors or promoting orthogonality among them. However, they overlook the fundamental target of model merging: the merged model performs as closely as possible to task-specific models on respective tasks. We find these methods inevitably discard task-specific information that, while causing conflicts, is crucial for performance. Based on our findings, we frame model merging as a constrained optimization problem (i.e., minimizing the gap between the merged model and individual models, subject to the constraint of retaining shared knowledge) and solve it via adaptive projective gradient descent. Specifically, we align the merged model with individual models by decomposing and reconstituting the loss function, alleviating conflicts through data-free optimization of task vectors. To retain shared knowledge, we optimize this objective by projecting gradients within a shared subspace spanning all tasks. Moreover, we view merging coefficients as adaptive learning rates and propose a task-aware, training-free strategy. Experiments show that our plug-and-play approach consistently outperforms previous methods, achieving state-of-the-art results across diverse architectures and tasks in both vision and NLP domains. Our code is available here.
@@ 1 Introduction
Fine-tuning pre-trained foundational models to address downstream tasks has become an effective paradigm (Muqeeth et al., 2024). However, the independent deployment of multiple fine-tuned models increases storage costs. While traditional multi-task learning (MTL) can mitigate these issues, they typically require concurrent training across multiple task-specific datasets, which incurs significant training overhead and potential privacy risks (Wei et al., 2024). Consequently, there is a growing interest in merging multiple expert models into a unified model without accessing their original data (Yang et al., 2024a; Huang et al., 2024). Model merging is performed directly at the parameter level and maintains only one final model during inference. In recent years, numerous pre-trained and fine-tuned checkpoints have been released on open-source communities like GitHub or Hugging Face, making it easy to obtain expert models from diverse domains. These rich model repositories underscore the value of model merging.
One popular approach, Task Arithmetic (TA) (Ilharco et al., 2023), combines task vectors through arithmetic operations for model merging. A major challenge is addressing conflicts that emerge when multiple task-specific models coexist within a single model. Ties-Merging (Yadav et al., 2023) proposes pruning redundant parameters, resolving sign conflicts, and merging sparse models, while AdaMerging (Yang et al., 2024c) applies test-time adaptation techniques to adjust merging coefficients in the weight space. Most recently, AWD (Xiong et al., 2024) finds that orthogonality among task vectors is key to model merging and introduces adaptive weight disentanglement to improve orthogonality. However, these methods overlook the fundamental requirement of model merging: ensuring the merged model performs comparably to task-specific models on their respective tasks.
@@
Based on our rethinking, we frame model merging as a constrained optimization problem (i.e., minimizing the gap between the merged model and individual models, subject to the constraint of retaining shared knowledge) and solve it via an adaptive projective gradient descent (doGe) method. Specifically, we measure the gap between the merged model and individual models in task-specific losses, and decompose it into a data-free objective using the first-order Taylor expansion. To alleviate conflicts, we introduce a modification vector $\Delta$ (i.e., redundant parameters) to each task vector. This data-free objective aims to achieve optimal average performance across multiple tasks by optimizing $\Delta$. For the modification vector, task vectors still compete to minimize the loss on their own tasks. Therefore, we construct a shared subspace based on all task vectors and optimize the problem within this subspace. The gradient of $\Delta$ can be divided into two components: one projected onto the shared subspace and the other orthogonal to it. We only take gradient steps in the direction orthogonal to the shared space, effectively constraining task vector optimization. As the former represents movements of parameters within the shared subspace, and the latter maintains shared knowledge while minimizing the gap for each task. Moreover, we determine task-aware, training-free merging coefficients based on the norm of task vectors to mitigate the dominance of any single task’s gradient influence.
We conduct experiments on diverse vision and NLP tasks, including classification and generation, using various fully fine-tuned and LoRA fine-tuned architectures. Our plug-and-play approach achieves up to 11.6% gains over TA and 5.8% over AdaMerging. Simple task-aware $\lambda$ provides a 2.8% performance boost. Furthermore, experiments on unseen tasks and out-of-distribution test sets demonstrate its generalization and robustness. Extensive ablation studies clarify the mechanisms of each component.
@@ 2 Related Work
Model merging (Crisostomi et al., 2024; Wang et al., 2024b; Daheim et al., 2024; Chen et al., 2024; Maldonado et al., 2024) eliminates the need for raw training data or expensive computations. It operates directly at the parameter level and consolidates multiple models into a single final model for inference. Existing model merging methods are categorized into two paradigms: pre-merging and during-merging (Yang et al., 2024a). Pre-merging methods aim to create favorable conditions for merging, such as using linearized fine-tuning to achieve weight disentanglement (Ortiz-Jimenez et al., 2023; Tang et al., 2024c).
@@
During-merging methods focus on developing techniques to merge given models and can be broadly categorized into data-free and test-time adaptation (TTA) approaches. TTA methods assume access to unlabeled test datasets and are often considered a form of transductive learning. For example, AdaMerging (Yang et al., 2024c) learns merging coefficients by minimizing entropy as a surrogate loss on test data, while Representation Surgery (Yang et al., 2024b) calibrates biases and aligns the merged model’s representations with those of the original task-specific models. In contrast, our approach designs a fully data-free objective to resolve task conflicts without relying on test data.
Data-free methods depend solely on the pre-trained and fine-tuned model weights for merging (Choi et al., 2024). Ties-Merging (Yadav et al., 2023) prunes redundant parameters by magnitude, resolves sign conflicts, and merges sparse models. Concrete Merging (Tang et al., 2023) adopts a meta-learning framework to learn a concrete mask that suppresses conflicting parameters. MAP (Li et al., 2025a) examines task vector magnitudes and leverages a second-order Taylor expansion to approximate loss-based metrics, providing a formal bound on the remainder term and using linear regression to estimate the Hessian.
Calculating the loss gap has been reflected in some studies: MetaGPT (Zhou et al., 2024) formally defines the loss difference and derives a closed-form solution for the merging coefficient $\lambda$. TATR (Sun et al., 2025) introduces the concept of knowledge conflict between tasks by modeling the loss gap as the product of gradients and task vectors. Other relevant works explore merging within subspaces. TSV (Gargiulo et al., 2024) aggregates task vectors using low-rank approximation and whitening to minimize interference, while KnOTS (Stoica et al., 2025) aligns representation spaces between LoRA models via SVD to enable compatible merging. These approaches, like ours, recognize the inherent low-rank structure of parameter updates and perform merging within subspaces. We focus on optimizing task vectors via gradient descent while constraining it within a shared subspace to retain shared knowledge.
Multi-task learning.
@@
Existing MTL research addresses the issue of negative transfer (Jiang et al., 2023) from two principal perspectives: architecture and optimization. From the architectural perspective, negative transfer is mitigated through strategies like modularization (Lu et al., 2024), sparsification (Sun et al., 2020), or soft sharing of the backbone. From an optimization perspective, it is widely recognized that tasks sharing similar underlying structures can benefit from being trained together. Gradient alignment methods (Yu et al., 2020; Shi et al., 2022) emphasize maintaining consistency in gradient directions or signs to resolve conflicts, which projects one task’s gradient onto the normal plane of another task’s gradient to reduce forgetting (Saha et al., 2021). Our approach enhances multi-task performance by aligning the merged model with each individual model and utilizing adaptive merging coefficients.
@@ 3.1 Preliminary
We begin with a pre-trained model $f$, parameterized by $\boldsymbol{\theta}_{0}$, which has been trained on a large-scale dataset. This model is paired with a collection of $n$ downstream tasks, denoted as $\{\mathcal{D}_{i}\}_{i=1}^{n}$. Subsequently, the pre-trained model $f$ is fine-tuned individually for each downstream task $\mathcal{D}_{i}$, resulting in a series of fine-tuned models, each with its unique parameters $\boldsymbol{\theta}_{i}$. To isolate task-specific information, we define the task vector as $\boldsymbol{\tau}_{i}=\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{0}$, a concept introduced by Ilharco et al. (2023). The set of these task vectors is represented as $\{\boldsymbol{\tau}_{i}\}_{i=1}^{n}$, enabling a focused analysis of the task-specific characteristics. Model merging aims to compose a multi-task model $\boldsymbol{\theta}^{*}$ to approximate the optimal solution:
@@ 3.2 Rethinking Model Merging for MTL
Resolving conflicts among tasks is a key challenge in model merging. Unlike MTL, which can mitigate conflicts during training with access to original data, model merging operates entirely in the parameter space. Existing methods mainly address conflicts by sparsely adjusting parameters, either by dropping conflicting parameters based on signs (Yadav et al., 2023) or importance scores (Du et al., 2024). Other methods promote orthogonality among task vectors, either by fine-tuning models in the tangent space (Ortiz-Jimenez et al., 2023) or directly optimizing task vectors (Xiong et al., 2024). While these methods alleviate conflicts, they inevitably discard task-specific information that contributes to conflicts, resulting in performance degradation. However, they overlook the fundamental target of model merging: the merged model performs as closely as possible to task-specific models on respective tasks. As shown in Fig. 1, increasing the task numbers leads to a continuous performance decline across methods. This is because more tasks result in increased negative transfer, causing the discard of valuable conflict-related task-specific knowledge. Therefore, we propose explicitly modeling the gap between the merged model $\boldsymbol{\theta}^{*}$ and individual models $\boldsymbol{\theta}_{i}$. This transforms conflict resolution into an optimization problem that can be solved using gradient descent.
@@
Figure 3: An illustration of element magnitudes in the task vector, inspired by (Shen et al., 2024). Best viewed when zoomed in.
@@
A critical observation is the importance of the merging coefficient $\lambda$. In methods like TA, a unified $\lambda$ is searched on the validation set. Ideally, $\lambda$ values should be task- and layer-specific. However, when dealing with a large number of tasks and layers, traditional methods such as grid search or combinatorial optimization search (Liu et al., 2020) become impractical. TTA methods require training $\lambda$ using unlabeled test data, which also presents limitations. A statistical analysis of task vector values reveals that tasks and layers exhibit different magnitudes (see Fig. 3). Modern adaptive optimizers (e.g., Adam) dynamically adjust learning rates based on gradient history, which is often more effective than a global learning rate. These optimizers suppress parameters with large gradients and reward those with small gradients, smoothing gradient fluctuations. Similarly, task vectors $\boldsymbol{\tau}_{i}$ represent cumulative gradients for each task, and $\lambda$ can be viewed as a learning rate balancing gradients across multiple tasks.
@@ 4 Methodology
$\min_{\boldsymbol{\theta}^{*}}$ $\mathcal{L}(\Delta;\lambda_{i},\boldsymbol{\tau}_{i}):=\sum_{i=1}^{n}\operatorname{Gap}(\boldsymbol{\theta}^{*},\boldsymbol{\theta}_{i}),$ (2)
$\,\mathrm{s.t.}$ $\operatorname{S_{\text{share}}}(\boldsymbol{\theta}^{*},\boldsymbol{\theta}_{0}+\lambda\sum_{i=1}^{n}\boldsymbol{\tau}_{i})=0.$
Here, $\Delta$ is initialized as a zero tensor with the same shape as the task vector. The function $\operatorname{Gap}(\cdot,\cdot)$ measures the distance between two sets of parameters, while $\operatorname{S}_{\text{share}}(\cdot,\cdot)$ denotes the distance within the shared subspace. Then, we solve it via adaptive projective gradient descent:
@@ 4.1 A Data-Free ̵Objective
Considering the fundamental target that the merged model should perform comparably to its respective task-specific model for each task, we follow Zhou et al. (2024) to define the objective for resolving model merging as:
@@
Similarly, calculating the gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{j}(\boldsymbol{\theta}_{0})$ of the pre-trained model for task $j$ requires access to data $\mathcal{D}_{j}$, which is typically unavailable. As an alternative, we approximate this gradient using the task vector $-\boldsymbol{\tau}_{j}$, since the task vector can be interpreted as an accumulation of gradients. Under the Neural Tangent Kernel assumption (i.e., fine-tuning occurs in a linear regime), $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{j}(\boldsymbol{\theta}_{0})$ can be estimated as $k\boldsymbol{\tau}_{j}$ with $k<0$. Here, $\boldsymbol{\tau}_{j}=\boldsymbol{\theta}_{T}-\boldsymbol{\theta}_{0}=-\sum_{t=1}^{T}\alpha_{t}\nabla_{\boldsymbol{\theta}_{t}}\mathcal{L}_{j}(\boldsymbol{\theta}_{t})$, where $\alpha_{t}$ is the learning rate and $T$ is the number of update steps. Given that parameters remain near $\boldsymbol{\theta}_{0}$, we have $\nabla_{\boldsymbol{\theta}_{t}}\mathcal{L}_{j}(\boldsymbol{\theta}_{t})=\nabla_{\boldsymbol{\theta}_{0}}\mathcal{L}_{j}(\boldsymbol{\theta}_{0})$. Thus, we obtain $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{j}(\boldsymbol{\theta}_{0})=-\boldsymbol{\tau}_{j}/\sum_{t=1}^{T}\alpha_{t}$. Consequently, the data-free objective can be approximated as:
@@ 4.2 Shared Subspace Optimization
We utilize Singular Value Decomposition (SVD) to extract the rank-$k$ subspace for the task vector. Specifically, the first $k$ singular vectors from the left singular matrix are selected as $\boldsymbol{B}_{j}$, forming an orthogonal basis that efficiently captures the primary information within the task-specific $\boldsymbol{\tau}_{j}$. Once the subspaces for all tasks are established, they are combined into a shared subspace $S_{\text{share}}=span\{[\boldsymbol{B}_{1},...,\boldsymbol{B}_{n}]\}$. However, $S_{\text{share}}$ includes overlapping singular vectors, indicating redundant parameters in the weight space across tasks. Such overlaps challenge the orthogonality requirement of basis vectors and lead to inaccuracies during projection onto the shared subspace. To mitigate this, we perform another SVD on $S_{\text{share}}$ to deduplicate it further, resulting in a refined $S_{\text{share}}$ that effectively preserves shared knowledge.
@@
Compared to optimizing $\Delta$ in the original parameter space, our approach explicitly constrains the gradient directions the optimizer can take. By taking gradient steps in the direction orthogonal to the shared subspace, we narrow the gap with the task-specific model. This effectively mitigates task conflicts while retaining shared knowledge.
@@
       Perform $\mathrm{SVD}$ on $\boldsymbol{\tau}_{i}$: $\boldsymbol{\tau}_{i}=U_{i}\Sigma_{i}V_{i}^{\top}$
@@
$S_{\text{share}}\leftarrow$ the first $k$ columns of $U$ from $\mathrm{SVD}\bigl{(}[\boldsymbol{B}_{1},\dots,\boldsymbol{B}_{n}]\bigr{)}$